UT May 05 2022 - Split Night - DEC / PEPSI-PETS



Night Info

  • Sunset 19:04
  • Nautical Twilight Ends 20:02
  • Astro Twilight Ends 20:35
  • Moon sets 23:04
  • Astro Twilight Starts 03:57
  • Nautical Twilight Starts 04:30
  • Sunrise 05:29

Staff

  • OSA: J.Williams
  • ISp: J.Schuh
  • Observer/ISA: A.Becker, O. Kuhn
  • Mtn Manager: R.Hansen

Night Plan

  • Starting Configuration: LBCs
    • Preparations:
      • updated lbcfpia.pro to allow z22off in response to dofpia, /askme (old version is lbcpfia_20220504_original.pro) -opk
      • made a copy of the tms_loop script which will send corrections when M1 is at the extrafocal position
        • cp tms_loop.pl tms_loop_x.pl
        • in tms_loop_x.pl, add "-x" after "-s 2" in the call to lbc_tms.py, like
          • system("${lbto::path::support}/MetrologySupport/lbc_tms.py -s 2 -x --tmspy '-vv --stdevpop 3 --stdevbnd 5 --mcp both'");
      • copied the TMS reference from 20220504 to telescope@robs:/tmp
  • LBC/TMS Engineering (primarily for LBC-B) - 20:35-
    • LBC Extra Focal images in equilibrium -IT 8608, IT 8592 - 1hr
    • Thermal Tracking - TMS Active - need thermal gradient - or create one with adjustment to mirror ventilation -1.5-2hr
  • LBC Science -1hr
    • TMS with channel dropping compensation turned on.
    • mid-z ULIRG imaging - 1hr
      • OT: /home/lbto/Service_Observing/2022A_D_E/MODS/Rothberg/midzUlirgs/midz_ULIRGs_V13.xml
        • LBC scripts are in LBC_Science folder, sub-folder for co-pointing scripts
        • rslaon_FF1106_Chip2_FOV.png finder for verification
        • Co-pointing finder FF1106_LBC_CoPointingFinder.png
      • Deep r-sloan binocular imaging of z~0.9 merger
      • r=23.1 (visible in 1 exposure)
      • 120sec exposures - 20pt dither - 20" and 25"
      • w/co-pointing and overheads should take 1hr total
  • Reconfigure for PEPSI 00:30
  • PEPSI Engineering time - 01:00-02:00hr (local)
    • Verify GCS hotspots
    • Observe flux standard to check throughput of new fibers
      • HR5191 RA 13hr Dec +49
  • PEPSI/PETS 12 - 02:00hr-04:40hr (local) -2hr40min

Poor Weather Backup

  • LUCI-Bino
    • Rothberg Local Mergers
      • OT: /home/lbto/Service_Observing/2022A_D_E/PoorWeather/Rothberg/LocalMergers/COBandhead_Mergers_V47.xml
        • 1 hour per target - 1hr15 with telluric
        • Not in order of priority - observe best placed target
        • NGC 4004 - DEC +27 - twilght-02hr
        • UGC 9829 - DEC -01 - 21:30-twilight
        • NGC 3597 - DEC -23 - twilight-22hr
    • Photometric Standards - any seeing, but must be clear
      • /home/lbto/Service_Observing/2022A_D_E/PoorWeather/LUCI_Phot_Standards
      • Select stars based on visibility, need to get stars at two different airmasses
  • LBCs
    • Photometric Standards - any seeing, but must be clear
      • OT File: /home/lbto/Service_Observing/2022A_D_E/LBC/New_PhotFields/New_Phot_Fields_Master_OT_V1.1.xml
  • MODS
    • Spetrophotometric Standards
      • Observe all modes/configurations - any seeing, but must be clear

Night Log

01:20 Preparations (see above):
  • updated lbcfpia.pro to allow z22off in response to dofpia, /askme (old version is lbcpfia_20220504_original.pro) -opk
    • made a copy of the tms_loop script which will send corrections when M1 is at the extrafocal position
      • cp tms_loop.pl tms_loop_x.pl
      • in tms_loop_x.pl, add "-x" after "-s 2" in the call to lbc_tms.py, like
        • system("${lbto::path::support}/MetrologySupport/lbc_tms.py -s 2 -x --tmspy '-vv --stdevpop 3 --stdevbnd 5 --mcp both'");
    • copied the TMS reference from 20220504 to telescope@robs:/tmp

01:20 Going to start TeamViewer - it tells me that I cannot because I have already the maximum number of concurrent sessions. I got around the problem by using a commercial license, however this does not help others unless that license can be made available to them.

02:05 Josh opened the enclosure

02:10 Running tms_loop.py. It is giving me some errors, both with --mcp both and --mcp none. I called Yang - the problem may be that the format of the reference has changed, so we cannot use the old reference today.

02:11 Slewed to a blank field for Uspec+F970N20 skyflats. Alex is taking skyflats.

02:33 Taking a new reference and using it. That works.

Pupil grid

02:36 Slewing to a field that is fairly high up to collimate and copoint. 75 deg elevation.
  • TMS update - 02:47
  • Stop TMS
  • dofpia
    • log file was not being written into /lbt/data/logs/lbc/fpia - corrected that
  • dofpia again
    • the estimated seeing is a bit high: 1.85 blue but 1.12 red. Blue is not doing any z11/z22 corrections because the seeing is too poor.
    • DIMM seeing is variable - 1.25 - 1.5".
  • setting the reference
  • taking the copointing exposures.
    • Sent the copointing OB - there was a large jump ~ -500 mic in group rotation on SX at 02:50:40. Group rotation is consistently the same on SX and the first image is in-focus. But now there has been a large positive jump in DX group rotation, from -0.09 to 0.72 mm. The first LBCR image was out of focus. This may be due to the fact that the new reference I just took was with the mirror at -0.8mm.
    • FWHM 3.8 Blue and 3.5 Red, though red is not in great focus.
  • Running dofpia again - the field only had a few good pupils (I made a mistake in telling the catalog what magnitude range to look for).
  • Running dofpia again on a slightly different field with more pupils, ~10.
  • Run copointing OB
    • set reference, this time wit the mirrors in focus, while taking the copointing images
  • copoint: offsets 3.2" ZB and 11.2" Red
  • Collect grid:
    • addsph.py -s 2
    • RB_fastextra4.ob
    • Step through (offsets) z11 = +/-1500nm, z22=+/200, +/400 and back to z11=z22=0
    • dofpia to finish
z11 z22 B R
1500 0 33320 33315
-1500 0 33549 33553
0 200 33829 33824
0 -200 34022 34018
0 -400 34216 34212
0 400 34435 34438
0 0 34649 34645
03:50 Completed the pupil grid, taking advantage of the fact that we were in pretty good equilibrium to start. But the DIMM seeing has degraded.
  • Copointing field has FWHM 5.5 Red and 5.4 Blue. The DIMM seeing is 1.5".

Thermal test

03:58 Seeing is marginal - but think we'll start collection of data for the thermal test.
  • dofpia, /askme (when z22off is given, it does not write the results to the log monitor - to the log too? I will need to check this - this is a good task for terrible conditions).
  • set the reference here, before slewing to the new field (mirrors are at -0.8mm)
  • ./tms_loop.pl
  • Slew to constant elevation field - now, wouldn't know you, the wind is from the north. It was from the SW at the start of the night. But it is only ~5m/s.
  • copointing: B 22.7" and R 19.4" from the rotator centers. Image quality - nice on red but a bit shabby on blue.
  • dofpia - there were large focus corrections on both sides. The "est seeing" is 1.45" B and 1.2" R
  • lbc_set_tms_offset.py -s 2
  • addsph.py -s 2
  • lbc_set_tms_offset.py -s 2 (again, just in case - what I did the last time)
  • note - the global offsets in the Zernikes z4 and z11 reflect two z4/z11 offsets, but this is because, after taking the pupil grid, I neglected to remove the corrections. We have collimated numerous times since then and have therefore absorbed the global offsets.

04:27 RB_fastextra4 series is starting
  • glass=4.5, ambient=4.45 on blue and glass = 4.35, 4.38 on red

04:35 Josh is changing the setpoint to 1.5C.

04:38 I am starting tms_loop_x.pl

05:16 Josh is going to drop the setpoint more, to -1.5C now. At this time glass=4.3, ambient 4.1 on SX and 4.084, 3.997 on DX
  • I have noticed spherical, but it was there from the start. The pupils do not seem to have changed.
  • We are just chasing the ambient temperature as it drops...

05:25 Josh is going to drive the mirror warmer, setting the setpoint to 6 C.

05:57 My lbto@robs session is frozen - user telescope's IDL session is taking 98% of the CPU. But I am not sure that is the culprit. I had run up another ds9 aside from RB_Science. I killed this ds9, 2 defunct ds9's and I suspended & restarted my x2go session and that helped.

06:07 We noticed that LBCR had stopped taking images. 054712 was the last. Rotator error.
  • Replaying the OB did not help
  • lbckill, lbcstart - that worked. 061221/061218 are the next B/R images.

06:26 The temperatures are not changing that much (glass=3.77, ambient 3.624 on SX and g=3.5=a on DX).
  • The seeing has also degraded. We'll stop here.

06:37 Stopping
  • remsph.py -s 2
  • dofpia - converged on the first iteration on the red side and needed some defocus. Estimated seeing 1.66, 1.54".
  • set the reference and started the active loop

Observing a high elevation field through transit

06:43 Slewing to a field that can be observed during meridian crossing at relatively high elevation (77 deg). The goal is to see whether the coma-buildup seen on 20220418 is reproducible.

06:42 Copointing. delta 8.6", -11.2" Collimation is a bit off on both sides.

06:53 dofpia --- Large corrections in z4, z11 and z22. but it converged in 2 iterations (seeing 1.5, 1.3")

06:57 Setting the reference, injected z4/z11 and then setting the reference again (to be sure there is nothing being missed). Started the tms_loop_x.pl

06:59 Starting the observations - 4-chip pupil images. 2 minutes before transit.

07:28 Stopped. dofpia (I did not remove the spherical z4/z11) The first iteration had all z5-z22 within convergence, only a z4 correction was needed. Ther had not been a buildup of coma as we saw on the LBCR images from 20220418. One difference, the rotator angle was 180 deg this time whereas it was 0 last time.

07:30 Handed over to PEPSI.

Summary

The seeing was marginal for the LBC engineering work, and the telescope was in too-good equilibrium. We collected the grid of pupil images at the start of the night, given that the seeing was expected to degrade and the temperature differential was pretty low. Then we moved on to try to collect pupil images as the mirror temperatures were driven up and down. Despite using a setpoint several degrees below and then several degrees above ambient, the mirrors changed temperature only very slowly, almost indiscernably, and the pupil images looked fairly constant throughout.

The field used for this test was a circumpolar field. Afterwards, and with seeing still marginal around 1.5", we spent the remaining time observing a high elevation field through transit, to see whether coma might develop, as it did on 20220418. The pupils did not show much evidence of coma.

The TMS code with channel dropping was used all night. John pointed out some issues on the DX side, but not on SX.

By collimating high and slewing low, and then collimating low and slewing to high elevation, we tested how well TMS holds collimation. In each case, it was pretty good although the first pair of images suggested that a dofpia run was needed.

I tested the modified lbcfpia code. It turns off sending z22 when the response to dofpia,/askme is "z22off". However, the FPIA logfile still records the original and not the actual z22, z4 and z11 values that are sent, so I will work on changing this. I have left this code in place (the old one is lbcfpia_20220505_original.pro).

-- BarryRothberg - 02 May 2022
I Attachment Action Size Date Who Comment
blue_pupil_grid_20220505.pngpng blue_pupil_grid_20220505.png manage 1 MB 05 May 2022 - 07:35 OlgaKuhn blue pupil images with z11=+/-1500, z22=+/-200, 400 and finally z11=z22=0.
red_pupil_grid_20220505.pngpng red_pupil_grid_20220505.png manage 1 MB 05 May 2022 - 07:33 OlgaKuhn red pupil images with z11=+/-1500, z22=+/-200, 400 and finally z11=z22=0.
Topic revision: r11 - 05 May 2022, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback